OpenStudy (goformit100):

Prove that the product of 4 consecutive natural numbers cannot be a perfect cube.

OpenStudy (amistre64):

n+n+1+n+2+n+3 = 4n + 3 a perfect cube (n+a)^3 = n^3 +3n^2a + 3na^2 + a^3 compare the two rewults

OpenStudy (amistre64):

oh, product ....

OpenStudy (amistre64):

you get n^4 + ....

OpenStudy (goformit100):

ok

OpenStudy (amistre64):

n * n+1 * n+2 * n+3 = n^4 + ... for n >1

OpenStudy (amistre64):

\[n^4+6n^3+11n^2+6n\]

OpenStudy (goformit100):

Ok Thank you sir

OpenStudy (amistre64):

we could possibly compare this with x^3 and see if they are equal at some n in N

OpenStudy (goformit100):

ok i'll do that

OpenStudy (amistre64):

\[y=n^4+6n^3+11n^2+6n~:~y=n^3\] \[n^4+6n^3+11n^2+6n=n^3\] \[n^4+5n^3+11n^2+6n=0\] \[n(n^3+5n^2+11n+6)=0\] \[n^3+5n^2+11n+6\] by descartes sign rule, this has no positive solutons

OpenStudy (goformit100):

ok